OpenStudy (butterflyhope):

Preform the indicated operation: -[5 - 2(3^2 - 2)]. Please show work. Thank you! :)

OpenStudy (campbell_st):

start with the innermost set of brackets 3^2 = 9 so its 9 - 2 = 7 then its \[-1 \times [5 - 2 \times (7)] \] inside the square brackets you need to do 2 x 7 then the subtraction... lastly multiply the answer by -1 hope it helps

OpenStudy (butterflyhope):

Where'd you get the -1? And yeah it helps, thanks :)

jimthompson5910 (jim_thompson5910):

the negative sign outside the brackets is the same as saying -1 times everything in the square brackets

OpenStudy (butterflyhope):

Ohhhh thanks :)

jimthompson5910 (jim_thompson5910):

you're welcome